OpenStudy (derpisthewordorisit):

MEDAL AND FAN 4 1/5 + b= 9 3/5

OpenStudy (tkhunny):

Have you considered changing your mixed numbers to improper fractions?

OpenStudy (derpisthewordorisit):

I tried, but I still couldn't do it.

OpenStudy (anonymous):

u could subtract 4 1/5 and 9 3/5

OpenStudy (derpisthewordorisit):

A. B=4 4/5 B. B=5 1/5 C. B= 5 2/5 D. B=13 4/5

OpenStudy (derpisthewordorisit):

Okay.

OpenStudy (tkhunny):

"I tried" is an insufficient response. What did you do? How did you do it? What were the results?

OpenStudy (derpisthewordorisit):

It's not, I'm doing it on a piece of paper right now. Thanks @Jade_Winters I think I got it now.

OpenStudy (tkhunny):

As noted by Jade_Winters, you could just subtract the mixed number, in this case. I would not consider that a generalized method. However, correct results don't care how you find them.

OpenStudy (derpisthewordorisit):

@Jade_Winters I got a different answer, and It's not on the answer choices.

OpenStudy (derpisthewordorisit):

Nevermind, I did it again and got it.

OpenStudy (derpisthewordorisit):

OHH I GET IT NOW THANKS!

OpenStudy (anonymous):

no prob